OpenStudy (anonymous):

Let f:R->R, where f(x)=3+6x-2x^3. a) determine the values of x for which the graph of f has positive gradient. b) find the values of x for which the graph of f has increasing gradient.

OpenStudy (anonymous):

in simpler language is this a) find where the derivative is positive?

OpenStudy (anonymous):

yep i understood that, but i wasnt sure what to do next.

OpenStudy (anonymous):

\[f'(x)=6-6x^2=6(1-x^2)=6(1+x)(1-x)\] a parabola that opens down. it will be positive between the zeros, namely on \[(-1,1)\]

OpenStudy (anonymous):

"increasing gradient" i take to mean the second derivative is positive. so \[f''(x)=-12x\] which is positive on \[(-\infty,0)\] and negative on \[(0,\infty)\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw